OpenStudy (anonymous):

The volume, V, of a tank varies directly with its height, h. A tank 12 feet high holds 300 cubic feet. Write an equation relating V and h. What is the slope of the relationship?

OpenStudy (anonymous):

V/ h = 300/ 12 -> slope V = ...

OpenStudy (anonymous):

I don't quite understand your explanation. May you explain it clearly ? :)

OpenStudy (anonymous):

Can you find V from the above ratio V/ h?

OpenStudy (anonymous):

300.

OpenStudy (anonymous):

No, the RATIO!

OpenStudy (anonymous):

Which means V depends on the variable h!

OpenStudy (anonymous):

means " Write an equation relating V and h"